Presentation is loading. Please wait.

Presentation is loading. Please wait.

LINAC3 INSTRUMENTATION STATUS F.Roncarolo – 18 Feb 2015 F.Roncarolo 18 Feb 20151.

Similar presentations


Presentation on theme: "LINAC3 INSTRUMENTATION STATUS F.Roncarolo – 18 Feb 2015 F.Roncarolo 18 Feb 20151."— Presentation transcript:

1 LINAC3 INSTRUMENTATION STATUS F.Roncarolo – 18 Feb 2015 F.Roncarolo 18 Feb 20151

2 Last BI Update was on Nov 19 th 2014 (https://indico.cern.ch/event/353640/contribution/2/material/slides/1.pdf) Open Issues: Pepper Pot Faraday Cups SEM Grids F.Roncarolo 18 Feb 20152

3 Pepper Pot Issue: cooling circuit vacuum leak Decision to have it inspected and repaired at Pantechnik Report received on Jan 15 th (document here)document here F.Roncarolo 18 Feb Results of tests after the detector reception at Pantechnik (Nov 2014): 1st Cycling test : 10 cycles at 5 bars during 15 min = OK 1st Long Term Stability test 24h at 5 bars = OK 2nd Long Term Stability test 15h at 5 bars = OK 2nd cycling test : 13 cycles at 5 bars during 15 min = OK 3rd Long Term Stability test (during the week end) = major leak after 49 hours and turn the leak detector in default mode (> mbars.l.s-1 ) A leak test has been performed on the water circuit, the leak is detected only when the pressure is increased over 2 bars. The water flange has been disconnected from the Tantalum screen and each side of the circuit has been tested. We detected a leak on both side but it was impossible to determine if the leak appeared on the bellow or the ceramic. Finally it was decided to build a new flange with a new water circuit (bellows, ceramics, fittings, solders). Result of tests with new cooling circuit (Jan 2015) : 10 cycles at 5 bars during 15 min = OK 1st Long Term Stability test 55h at 5 bars = OK 2nd Long Term Stability test 17h at 5 bars = OK Leak rate always stable at ~2, mbars.l.s-1

4 Pepper Pot Detector was sent back to CERN Reception on Jan 14 th Decision: TE-VSC to validate the system before instalaltion  Installation not possible by end of LINAC3 YETS (Jan 16 th ) TE-VSC (Alice) completed long term pressure cycles this week Looks OK  Install detector at next LINAC3 stop ? F.Roncarolo 18 Feb 20154

5 Faraday Cup(s) Old Issue: negative peaks (from shower of secondary electrons?) BI-ML looking into new design (no dates yet) New Issue Sparking with Argon Related to inspection December 2014? We re-put signal and bias cables as they were before If you believe there is a bad wiring, we need to open Do we have expected beam distributions (Ion species and transverse distributions) ? With this we could simulate Old and New issue F.Roncarolo 18 Feb 20155

6 SEM grids Issue: grid frame sticking into beam pipe aperture when in parking position  Charge collection / beam kick  Potential heating / damage After BI-ML (C.Vuitton) inspection (December 2014): 1. Changer le vérin FESTO course 100mm par un vérin FESTO DNC PPV (amortissement) course 125mm pour gagner 12 à 15mm de course direction « garage ». Remarques: pas nécessaire de changer le vérin d’équilibre AIRAX (course 120mm) ni les 2 soufflets. 2. Réduire l’ouverture du tank en installant une bague avec diam de passage faisceaux de 80mm. Actions : acheter le vérin Festo et fabriquer une ou 2 bagues de réductions. demander un arrêt de 1 jour : retirer le moniteur du LINAC 3. échanger le vérin. réglage des courses. (Nécessité de mettre le moniteur sur un banc d’alignement.) Installer la bague de réduction (diam 80mm) dans le tank. Remonter le moniteur. For the moment modification foreseen for ITL.MSG03, procurment of pieces launched for lab test validation If all ok: LINAC3 in 3 months F.Roncarolo 18 Feb 20156

7 SEM Grids – New requests (from BE-ABP, G.Bellodi) Can we change the number of wires and the wire ITL and ITM SEM grids? Electronics: easy to pass from 25 to 32 wires per plane Mechanics: BI-ML investigating, should be possible, need ~3 Months (re-design ceramic PCB wiring) Can we read-out L3 SEM-grids signal with time resolving along beam pulse? YES: electronics and FESA class already sampling ADC at 4us, just need to modify OP GUI F.Roncarolo 18 Feb 20157

8 SPARES F.Roncarolo 18 Feb 20158

9 New Pepper Linac3 Purchased from Pantechnik F.Roncarolo 18 Feb New diagnostics tank hosting Slit Pepper Pot FC Pepper Pot

10 F.Roncarolo 18 Feb

11 New Pepper Linac3 F.Roncarolo 18 Feb Here in the parking position BEAM Pepper pot plate MCP, PHOSPHOR Mirror Camera


Download ppt "LINAC3 INSTRUMENTATION STATUS F.Roncarolo – 18 Feb 2015 F.Roncarolo 18 Feb 20151."

Similar presentations


Ads by Google